# A system of NLS arising in optical material without Galilean symmetry

### v1

@@ 3 High frequency limit
$0=$ $N_{1,0}\left(\tilde{u}^{\lambda_{\omega}}_{\omega},\tilde{v}^{\lambda_{\omega}}_{\omega}\right)$
$=$ $\frac{\lambda_{\omega}^{N-2}}{4-N}\left(N(1-\lambda_{\omega}^{2})\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)+[N-2-2\lambda_{\omega}^{2}]\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})\right).$
@@
$\lambda_{\omega}^{N-2}S_{1,0}\left(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}}\right)$
$=$ $\frac{1}{2}\int\left(|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}\right)+\frac{(\lambda_{\omega})^{2}}{2}\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)-(\lambda_{\omega})^{2}D(\tilde{u}_{\omega},\tilde{v}_{\omega})$
$=$ $\frac{1}{2}\int\left(|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}\right)+\lambda_{\omega}^{2}\frac{1}{4}\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)$
@@
$=$ $\left(\frac{1}{2}-\frac{\lambda_{\omega}^{2}}{4}\right)\int\left(|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}\right)+\lambda_{\omega}^{2}\frac{1}{4}\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)-\lambda_{\omega}^{2}\frac{1}{4}\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})$
$=$ $\frac{1}{2}Q_{1,\frac{c}{\sqrt{w}}}(\tilde{u}_{\omega},\tilde{v}_{\omega})+\frac{1}{4}(1-\lambda_{\omega}^{2})\int\left(|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}\right)$
$+\frac{1}{4}(1-\lambda_{\omega}^{2})\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)-\frac{1}{4}(1+\lambda_{\omega}^{2})\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})$
[…]
where $Q_{1,\frac{c}{\sqrt{\omega}}}$ is defined by (2.1) and in the last step, we used the Lemma 3.3 and (3.11). This means that
$\mu_{1,0}\leq\mu_{1,\frac{c}{\sqrt{\omega}}}+o(1),~{}~{}\text{as}~{}~{}\omega\to+\infty.$ (3.12)
@@
Up to a subsequence, $(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{u}_{\omega}^{\lambda_{\omega}})\to(\tilde{u}_{\infty},\tilde{v}_{\infty})$ strongly in $H^{1}$ as $\omega\rightarrow+\infty$. Moreover, $(\tilde{u}_{\infty},\tilde{v}_{\infty})$ is a minimizer sequence of $\mu_{1,0}$.
@@
$N_{1,0}(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{u}_{\omega}^{\lambda_{\omega}})=0$
@@
$S_{1,0}(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{u}_{\omega}^{\lambda_{\omega}})\to\mu_{1,0}~{}~{}\text{as}~{}~{}\omega\to\infty.$
This means that $\left(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}}\right)$ is the minimizing sequence of (2.3). By the definition of $(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{u}_{\omega}^{\lambda_{\omega}})$ (see (3.9)) and Lemma 3.3, we get $(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{u}_{\omega}^{\lambda_{\omega}})$ is also uniformly bounded in $H^{1}(\mathbb{R}^{N})\times H^{1}(\mathbb{R}^{N})$.
By similar argument as Lemma 2.6 and 2.9, there exist $\{y_{\omega}\}$, $(\tilde{u}_{\infty},\tilde{v}_{\infty})\in H^{1}(\mathbb{R}^{N})\times H^{1}(\mathbb{R}^{N})\backslash\{(0,0)\}$, and a subsequence $\tau_{y_{\omega}}\left(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}}\right)$ (still denoted by $\tau_{y_{\omega}}\left(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}}\right)$ such that
@@
$Q_{1,0}\tau_{y_{\omega}}(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}})-Q_{1,0}\left(\tau_{y_{\omega}}(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}})-(\tilde{u}_{\infty},\tilde{v}_{\infty})\right)\to Q_{1,0}(\tilde{u}_{\infty},\tilde{v}_{\infty}).$
Since $Q_{1,0}(\tilde{u}_{\infty},\tilde{v}_{\infty})>0$, up to a subsequence, we have
$\frac{1}{2}\lim_{\omega\to\infty}Q_{1,0}\left(\tau_{y_{\omega}}(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}})-(\tilde{u}_{\infty},\tilde{v}_{\infty})\right)<\frac{1}{2}\lim_{\omega\to\infty}Q_{1,0}(\tau_{\omega}(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}}))=\mu_{1,0}.$
@@
$\mu_{1,0}\leq\frac{1}{2}Q_{1,0}(\tilde{u}_{\infty},\tilde{v}_{\infty})\leq\frac{1}{2}\lim_{\omega\to\infty}Q_{1,0}(\tau_{\omega}(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}}))=\mu_{1,0}.$
@@
$Q_{1,0}\left(\tau_{y_{\omega}}(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}})-(\tilde{u}_{\infty},\tilde{v}_{\infty})\right)\to 0,$
@@ 4 Global existence
$A^{+}_{\omega,c}=\{(u,v)\in H^{1}(\mathbb{R}^{N})\times H^{1}(\mathbb{R}^{N})~{}:~{}S_{\omega,c}(u,v)\leq\mu_{\omega,c},~{}N_{\omega,c}(u,v)\geq 0\}.$ (4.1)
Now we show that $A^{+}_{\omega,c}$ is an invariant set under the flow.
@@
Assume (1.11). Then the set $A^{+}_{\omega,c}$ is invariant under the flow of (1.1).
@@
By the similar argument as [9, 12], we can easily obtain this result. Here we omit it. ∎
@@
Assume the condition (1.11) holds. If the initial data $(u_{0},v_{0})\in A^{+}_{\omega,c}$, then the $H^{1}$ solution $(u(t),v(t))$ of (1.1) exists globally in time and
@@
Combining Lemma 4.1 and the definitions of $S_{\omega,c}$, $Q_{\omega,c}$ and $N_{\omega,c}$, we have
$\mu_{\omega,c}\geq$ $S_{\omega,c}(u(t),v(t))$
$=$ $\frac{1}{2}Q_{\omega,c}(u(t),v(t))+\frac{1}{4}N_{\omega,c}(u(t),v(t))$
$\geq$ $\frac{1}{2}Q_{\omega,c}(u(t),v(t))$
@@
$\leq$ $C\mu_{\omega,c}+CM_{3\gamma}(u_{0},v_{0}).$
@@
To prove (i). Let $A_{0}$ be chosen later. We claim that if $(u_{0},v_{0})\in H^{1}(\mathbb{R}^{2})$ satisfies $\|v_{0}\|_{L^{2}}^{2}<A_{0}$ and $0<\epsilon\ll\frac{1}{|c|^{2}}$ , then we claim that
$S_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c}\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)<\mu_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c},$ (4.2)
$N_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c}\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)\geq 0.$ (4.3)
@@
$\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)\in A^{+}_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c}$
@@
Now we prove the claim. Indeed, by the definition of $Q_{\omega,c}$, we have
$Q_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c}(u,v)$
@@
$S_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c}\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)$
$=$ $Q_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c}\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)-D\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)$
@@
$\leq$ $|c|^{2}\mu_{\frac{1+\epsilon}{4},\frac{c}{|c|}}.$
@@
$\leq$ $|c|^{2}\left(\mu_{\frac{1+\epsilon}{4},\frac{c}{|c|}}-\frac{\gamma(3(1+\epsilon)-\gamma)}{8}\|v_{0}\|_{L^{2}}^{2}\right).$ (4.4)
@@
$A_{0}=\frac{8}{\gamma(3(1+\epsilon)-\gamma)}\mu_{\frac{1+\epsilon}{4},\frac{c}{|c|}}.$
Then, by $\gamma<3$ and Lemma 2.5, we get $A_{0}>0$. Let $\|v_{0}\|_{L^{2}}^{2}<A_{0}$, then
$|c|^{2}\left(\mu_{\frac{1+\epsilon}{4},\frac{c}{|c|}}-\frac{\gamma(3(1+\epsilon)-\gamma)}{8}\|v_{0}\|_{L^{2}}^{2}\right)\to\infty~{}~{}\text{as}~{}~{}|c|\to\infty.$
@@
From (4.5), we deduce that the left-hand side of (4) is bounded as $|c|$ is sufficiently large, since $0<\epsilon\ll\frac{1}{|c|^{2}}$. Therefore, (4) holds for $|c|$ large enough, and so does (4.2). In addition, by (4.5), we obtain
$N_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c}\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)$
$=$ $2Q_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c}\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)-4D\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)$
@@
$N_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c}\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)\geq 0$
@@
To prove (ii). Let $0<\epsilon\ll\frac{1}{|c|^{2}}$ and
$B_{0}=\frac{24}{(1+\epsilon)\gamma-3}\mu_{\frac{(1+\epsilon)\gamma}{12},\frac{c}{|c|}}.$

### v2

@@ 3 High frequency limit
Case 3. Assume that $\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})>0$ and $\alpha-3\gamma<0$. If $N_{1,0,0}(\tilde{u}_{\omega},\tilde{v}_{\omega})<0$, then by the similar argument as Case 1, there exists $\lambda_{\omega}\in(0,1)$ such that $N_{1,0,0}(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}})=0$; if $N_{1,0,0}(\tilde{u}_{\omega},\tilde{v}_{\omega})>0$, then by the similar argument as Case 2, there exists $\lambda_{\omega}\in(1,\infty)$ is uniformly bounded, such that $N_{1,0,0}(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}})=0$.
Case 4. Assume that $\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})>0$ and $\alpha-3\gamma>0$. By the same argument as Case 3, we can obtain the similar result. Here we omit it.
@@
$0=$ $N_{1,0,0}\left(\tilde{u}^{\lambda_{\omega}}_{\omega},\tilde{v}^{\lambda_{\omega}}_{\omega}\right)$
$=$ $\frac{\lambda_{\omega}^{N-2}}{4-N}\Bigg{(}N(1-\lambda_{\omega}^{2})\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)+[N-2-2\lambda_{\omega}^{2}]\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})\Bigg{)}.$
@@
$\lambda_{\omega}^{N-2}S_{1,0,0}\left(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}}\right)$
$=$ $\frac{1}{2}\int\left(|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}\right)+\frac{\lambda_{\omega}^{2}}{2}\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)-\lambda_{\omega}^{2}D(\tilde{u}_{\omega},\tilde{v}_{\omega})$
$=$ $\frac{1}{2}\int\left(|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}\right)+\lambda_{\omega}^{2}\frac{1}{4}\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)-\frac{\lambda_{\omega}^{2}(\alpha-3\gamma)}{4\omega}\|\tilde{v}_{\omega}\|_{L^{2}}^{2}$
@@
$=$ $\left(\frac{1}{2}-\frac{\lambda_{\omega}^{2}}{4}\right)\int\left(|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}\right)+\lambda_{\omega}^{2}\frac{1}{4}\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)$
$-\frac{\lambda_{\omega}^{2}(\alpha-3\gamma)}{4\omega}\|\tilde{v}_{\omega}\|_{L^{2}}^{2}-\lambda_{\omega}^{2}\frac{1}{4}\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})$
$=$ $\frac{1}{2}\mathcal{Q}_{1,\frac{c}{\sqrt{w}},\frac{\alpha-3\gamma}{\omega}}(\tilde{u}_{\omega},\tilde{v}_{\omega})+\frac{1}{4}(1-\lambda_{\omega}^{2})\int\left(|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}\right)+\frac{(1-\lambda_{\omega}^{2})(\alpha-3\gamma)}{4\omega}\|\tilde{v}_{\omega}\|_{L^{2}}^{2}$
$+\frac{1}{4}(\lambda_{\omega}^{2}-1)\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)-\frac{1}{4}(1+\lambda_{\omega}^{2})\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})$
$=$ $\mu_{1,\frac{c}{\sqrt{\omega}},\frac{\alpha-3\gamma}{\omega}}+o(1),~{}~{}\text{as}~{}~{}\omega\to+\infty.$
where $\mathcal{Q}_{1,\frac{c}{\sqrt{\omega}},\frac{\alpha-3\gamma}{\omega}}$ is defined by (2.1) and in the last step, we also used the Lemma 3.3, (3.17) and (3.14). This means that
$\mu_{1,0,0}=\mu_{1,\frac{c}{\sqrt{\omega}},\frac{\alpha-3\gamma}{\omega}}+o(1),~{}~{}\text{as}~{}~{}\omega\to+\infty.$ (3.18)
@@
Up to a subsequence, $(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{u}_{\omega}^{\lambda_{\omega}})\to(\tilde{u}_{\infty},\tilde{v}_{\infty})$ strongly in $H^{1}\times H^{1}$ as $\omega\rightarrow+\infty$. Moreover, $(\tilde{u}_{\infty},\tilde{v}_{\infty})$ is a minimizer sequence of $\mu_{1,0,0}$.
@@
$N_{1,0,0}(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{u}_{\omega}^{\lambda_{\omega}})=0$
@@
$S_{1,0,0}(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{u}_{\omega}^{\lambda_{\omega}})\to\mu_{1,0,0}~{}~{}\text{as}~{}~{}\omega\to\infty.$
This means that $\left(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}}\right)$ is the minimizing sequence of (2.3) with $c=0$ and $\alpha=3\gamma$. By the definition of $(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{u}_{\omega}^{\lambda_{\omega}})$ (see (3.15)) and Lemma 3.3, we get $(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{u}_{\omega}^{\lambda_{\omega}})$ is also uniformly bounded in $H^{1}(\mathbb{R}^{N})\times H^{1}(\mathbb{R}^{N})$.
By similar argument as Lemma 2.5 and 2.8, there exist $\{y_{\omega}\}$, $(\tilde{u}_{\infty},\tilde{v}_{\infty})\in H^{1}(\mathbb{R}^{N})\times H^{1}(\mathbb{R}^{N})\backslash\{(0,0)\}$, and a subsequence $\tau_{y_{\omega}}\left(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}}\right)$ (still denoted by $\tau_{y_{\omega}}\left(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}}\right)$ such that
@@
$\mathcal{Q}_{1,0,0}\tau_{y_{\omega}}(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}})-\mathcal{Q}_{1,0,0}\left(\tau_{y_{\omega}}(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}})-(\tilde{u}_{\infty},\tilde{v}_{\infty})\right)\to\mathcal{Q}_{1,0,0}(\tilde{u}_{\infty},\tilde{v}_{\infty}).$
Since $\mathcal{Q}_{1,0,0}(\tilde{u}_{\infty},\tilde{v}_{\infty})>0$, up to a subsequence, we have
$\frac{1}{2}\lim_{\omega\to\infty}\mathcal{Q}_{1,0,0}\left(\tau_{y_{\omega}}(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}})-(\tilde{u}_{\infty},\tilde{v}_{\infty})\right)<\frac{1}{2}\lim_{\omega\to\infty}\mathcal{Q}_{1,0,0}(\tau_{\omega}(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}}))=\mu_{1,0,0}.$
@@
$\mu_{1,0,0}\leq\frac{1}{2}\mathcal{Q}_{1,0,0}(\tilde{u}_{\infty},\tilde{v}_{\infty})\leq\frac{1}{2}\lim_{\omega\to\infty}\mathcal{Q}_{1,0,0}(\tau_{\omega}(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}}))=\mu_{1,0,0}.$
@@
$\mathcal{Q}_{1,0,0}\left(\tau_{y_{\omega}}(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}})-(\tilde{u}_{\infty},\tilde{v}_{\infty})\right)\to 0,$
@@ 4 Global existence
$A^{+}_{\omega,c,\alpha-3\gamma}=\{(u,v)\in H^{1}(\mathbb{R}^{N})\times H^{1}(\mathbb{R}^{N})~{}:~{}S_{\omega,c,\alpha-3\gamma}(u,v)\leq\mu_{\omega,c,\alpha-3\gamma},~{}N_{\omega,c,\alpha-3\gamma}(u,v)\geq 0\}.$ (4.1)
Obviously, $A^{+}_{\omega,c,\alpha-3\gamma}$ is nonempty.
Now we show that $A^{+}_{\omega,c,\alpha-3\gamma}$ is an invariant set under the flow.
@@
Assume (1.10). Then the set $A^{+}_{\omega,c,\alpha-3\gamma}$ is invariant under the flow of (1.1).
@@
Let $(u_{0},v_{0})\in A^{+}_{\omega,c,\alpha-3\gamma}$. It is obvious that $S_{\omega,c,\alpha-3\gamma}(u(t),v(t))\leq\mu_{\omega,c,\alpha-3\gamma}$ for all $t\in I$, where $I$ is the maximal existence interval of $H^{1}$-solution, since the corresponding mass, energy and momentum are conserved.
Now we show that $N_{\omega,c,\alpha-3\gamma}(u(t),v(t))\geq 0$ for all $t\in I$. If not, there exist $t_{1},t_{2}\in I$ such that $N_{\omega,c,\alpha-3\gamma}(u(t_{1}),v(t_{1}))<0$ and $N_{\omega,c,\alpha-3\gamma}(u(t_{2}),v(t_{2}))=0$. By the uniqueness of Cauchy problem for (1.1), we have $(u(t_{2}),v(t_{2}))\neq(0,0)$. Moreover, from $S_{\omega,c,\alpha-3\gamma}(u(t),v(t))\leq\mu_{\omega,c,\alpha-3\gamma}$, we obtain $(u(t_{2}),v(t_{2}))\in a_{\omega,c,\alpha-3\gamma}\subset\mathcal{G}_{\omega,c,\alpha-3\gamma}$. This yields that
$(u(t),v(t))=\left(e^{i\omega(t-t_{2})}u(t_{2},x-c(t-t_{2})),e^{3i\omega(t-t_{2})}v(t_{2},x-c(t-t_{2}))\right)$
for all $t\in\mathbb{R}$. In particular, $N_{\omega,c,\alpha-3\gamma}(u(t),v(t))=0$ for all $t\in\mathbb{R}$, which contradicts $N_{\omega,c,\alpha-3\gamma}(u(t_{1}),v(t_{1}))<0$.
By similar argument as above, we can prove the case $A^{-}_{\omega,c,\alpha-3\gamma}$. Now we complete the proof of this lemma. ∎
@@
Assume the condition (1.10) holds. If the initial data $(u_{0},v_{0})\in A^{+}_{\omega,c,\alpha-3\gamma}$, then the $H^{1}$ solution $(u(t),v(t))$ of (1.1) exists globally in time and
@@
Combining Lemma 4.1 and the definitions of $S_{\omega,c,\alpha-3\gamma}$, $\mathcal{Q}_{\omega,c,\alpha-3\gamma}$ and $N_{\omega,c,\alpha-3\gamma}$, we have
$\mu_{\omega,c,\alpha-3\gamma}\geq$ $S_{\omega,c,\alpha-3\gamma}(u(t),v(t))$
$=$ $\frac{1}{2}\mathcal{Q}_{\omega,c,\alpha-3\gamma}(u(t),v(t))+\frac{1}{4}N_{\omega,c,\alpha-3\gamma}(u(t),v(t))$
$\geq$ $\frac{1}{2}\mathcal{Q}_{\omega,c,\alpha-3\gamma}(u(t),v(t))$
@@
$\leq$ $C\mu_{\omega,c,\alpha-3\gamma}+CM_{3\gamma}(u_{0},v_{0}).$
@@
Remark 4.3.
We note that if $\mu_{\omega,c,\alpha-3\gamma}$ is independent on the angle of $c$. Indeed, let $c_{1},c_{2}\in\mathbb{R}^{N}$ satisfy $|c_{1}|=|c_{2}|$. Then there exists an orthogonal matrix $R$ such that $Rc_{2}=c_{1}$. Let $(u_{n},v_{n})$ be a minimizing sequence for $\mu_{\omega,c_{1},\alpha-3\gamma}$, i.e., $(u_{n},v_{n})\in N_{\omega,c_{1},\alpha-3\gamma}$ and $S_{\omega,c_{1},\alpha-3\gamma}(u_{n},v_{n})\to\mu_{\omega,c_{1},\alpha-3\gamma}$. Let $(w_{n}(x),z_{n}(x))=(u_{n}(Rx),v_{n}(Rx))$, then
$N_{\omega,c_{2},\alpha-3\gamma}(w_{n},z_{n})=N_{\omega,c_{1},\alpha-3\gamma}(u_{n},v_{n})=0,$
$S_{\omega,c_{2},\alpha-3\gamma}(w_{n},z_{n})=S_{\omega,c_{1},\alpha-3\gamma}(u_{n},v_{n})\to\mu_{\omega,c_{1},\alpha-3\gamma}.$
This implies that $\mu_{\omega,c_{2},\alpha-3\gamma}\leq\mu_{\omega,c_{1},\alpha-3\gamma}$. Similarly, the inverse inequality holds. In particular, when $\alpha=3\gamma$, $\mu_{\omega,\frac{c}{|c|},0}$ is constant independent of $|c|$.
@@
To prove (i). Let $A_{0}$ be chosen later. We claim that if $(u_{0},v_{0})\in H^{1}(\mathbb{R}^{2})$ satisfies $\|v_{0}\|_{L^{2}}^{2}<A_{0}$ and $0<\epsilon\leq\frac{1}{|c|^{2}}$ , then we claim that
$S_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c,0}\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)<\mu_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c,0},$ (4.2)
$N_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c,0}\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)\geq 0.$ (4.3)
@@
$\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)\in A^{+}_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c,0}$
@@
Now we prove the claim. Indeed, by the definition of $\mathcal{Q}_{\omega,c,0}$, we have
$\mathcal{Q}_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c,0}(u,v)$
@@
$S_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c,0}\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)$
$=$ $\mathcal{Q}_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c,0}\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)-D\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)$
@@
$\leq$ $|c|^{2}\mu_{\frac{1+\epsilon}{4},\frac{c}{|c|},0}.$
@@
$\leq$ $|c|^{2}\left(\mu_{\frac{1+\epsilon}{4},\frac{c}{|c|},0}-\left(\frac{\gamma(3(1+\epsilon)-\gamma)}{8}\right)\|v_{0}\|_{L^{2}}^{2}\right).$ (4.4)
@@
$A_{0}=\frac{8}{\gamma(3(1+\epsilon)-\gamma)}\mu_{\frac{1+\epsilon}{4},\frac{c}{|c|},0}.$
From Lemma 3.2 and Remark 4.3, $\mu_{\frac{1+\epsilon}{4},\frac{c}{|c|},0}$ is independent on $c$. Then, by $\gamma<3$ and Lemma 2.4, we get $A_{0}>0$. Let $\|v_{0}\|_{L^{2}}^{2}<A_{0}$, then
$|c|^{2}\left(\mu_{\frac{1+\epsilon}{4},\frac{c}{|c|},0}-\left(\frac{\gamma(3(1+\epsilon)-\gamma)}{8}\right)\|v_{0}\|_{L^{2}}^{2}\right)\to 0,~{}~{}\text{as}~{}~{}|c|\to\infty.$
@@
From (4.5), we deduce that the left-hand side of (4) is bounded as $|c|$ is sufficiently large, since $0<\epsilon\leq\frac{1}{|c|^{2}}$. Therefore, (4) holds for $|c|$ large enough, and so does (4.2).
Again, by (4.5), we obtain
$N_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c,0}\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)$
$=$ $2\mathcal{Q}_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c,0}\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)-4D\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)$
@@
$N_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c,0}\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)\geq 0$
@@
To prove (ii). Let $0<\epsilon\leq\frac{1}{|c|^{2}}$ and
$B_{0}=\frac{24}{(1+\epsilon)\gamma-3}\mu_{\frac{(1+\epsilon)\gamma}{12},\frac{c}{|c|},0}.$